OpenStudy (anonymous):

Solve for x: -3|2x + 6| = -12 x = 1 and x = 5 x = -1 and x = -5 x = -9 and x = 3 No solution

OpenStudy (anonymous):

fist you have to divide -3 from both sides. then you get \[\left| 2x+6 \right|=4\] 2x+6=4 2x=-2 x=-1 2x+6=-4 2x=-10 x=-5 B would be the correct answer. -1,-5

OpenStudy (anonymous):

thanks